# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v2

@@ IV-A Experiment Design
TABLE II: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I.
@@
IV-A1 Simulation
The goal of our simulation experiments was to empirically test several properties of our algorithm under highly controlled settings. To this end, we used simulated T1W and T2W images based on BrainWeb phantoms [44]. BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing the reconstruction, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling at various accelerations. We analyzed the effect of the content/style model’s disentanglement level and its content capacity on the reconstruction quality, the convergence of the PnP-MUNIT algorithm, and the effectiveness, robustness, and tuning of the CR module.
IV-A2 Benchmark on NYU DICOM dataset
In the NYU benchmark, we compared PnP-MUNIT against end-to-end reconstruction methods. The question we sought to answer was whether or not PnP-MUNIT, which requires only image-domain training data, can outperform methods that additionally require k-space training data. The NYU brain dataset [45, 46] includes raw data and DICOM scans of 4 contrasts – T1W with and without gadolinium agent, T2W, and FLAIR. Following [41], a paired DICOM subset of 327 subjects was obtained based on T2W and non-gadolinium T1W scans. All T1W scans were rigid-registered with the corresponding T2W scans. Single-coil T2W k-space data was simulated via Fourier transform with 1D Cartesian random sampling at $R\in\{2,3,4,5\}$ and added Gaussian noise of $\sigma=0.01\mathrm{max}(x_{2}^{*})$. The dataset was split into 4 subsets – model-train (200), model-val (27), recon-val (50), and recon-test (50). Images from the former two splits were used to pre-train, fine-tune, and validate the MUNIT model. On the other hand, end-to-end models were trained and validated on the aligned T1W, the simulated T2W k-space, and the T2W ground-truth from these two splits. The recon-val split was used to tune the PnP-MUNIT algorithm, and recon-test was the held-out test set. We benchmarked PnP-MUNIT against one well-known single-contrast network – MoDL [4] – and two recent multi-contrast networks – MTrans [32] and MC-VarNet [15].
IV-A3 Benchmark and ablation on LUMC multi-coil dataset
@@
In our benchmark on clinical multi-coil data, our goal was to test PnP-MUNIT on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance. Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table I shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
@@
As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate MUNIT. The MUNIT models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines. An overview of the data split is shown in Table II. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
@@
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-MUNIT with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) [7]. As an unguided plug-and-play baseline, we used PnP-CNN [23], and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used (PROSIT) [39], which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound for PnP-MUNIT representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$. In both NYU and LUMC benchmarks, we used 3 perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality [47]. All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
IV-A4 Radiological evaluation
@@
Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-MUNIT reconstructions. The study was conducted on a small sample of the LUMC-TRA recon-test set. With the help of a junior radiologist, we selected 3 cases with brain metastases. The evaluation comprised two parts, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale [48] – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. A diagnostic-quality reconstruction was defined as one that showed at least 90% of the metastases and scored at least “fairly diagnostic quality” on all other criteria. [^2] In clinical practice, while the remaining 10% of the metastases may not be detected in the reconstructed T2W scan, they would be prominently visible in the corresponding gadolinium-enhanced T1W scan, and hence would not go undetected. We evaluated PnP-MUNIT and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to a senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
IV-B Implementation Details
@@
We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of given number of channels. We used 2 or 4 channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal. To stabilize GAN training, we used spectral normalization [49]. For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets. In the benchmarks, the CR module parameter $\gamma$ of PnP-MUNIT was tuned per acceleration for each dataset and the number of iterations was set to 200.
@@
In the simulations, all MUNIT models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section V-A. In both NYU and LUMC benchmarks, the MUNIT models were pre-trained for 400k iterations and fine-tuned for 50k iterations. All software was implemented in PyTorch, and image registration was performed using Elastix [50]. All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
V Results
@@
TABLE III: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than L1-wavelet CS is marked as .
@@
TABLE IV: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than L1-wavelet CS is marked as .
@@
RefRes-1 RefRes-2 RefRes-4 RefRes-8
ContentRes-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
ContentRes-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
ContentRes-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
Figure 3: (a) Convergence of PnP-MUNIT and the baseline versions at $R$=4. (b) Evolution of the reconstruction shown with its error map.
@@
V-A Experiments on Simulated MR Datasets
V-A1 Perturbation of disentanglement strength
PnP-MUNIT relies heavily on a sufficient disentanglement of content and style from the images. Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq. (6)) of MUNIT on the PnP-MUNIT reconstruction quality. Table III shows the reconstruction PSNR for different weight values. We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$. At lower values, poor disentanglement of content and style in the model resulted in poor PnP-MUNIT reconstruction quality. This was expected since PnP-MUNIT relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels. At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
V-A2 Content resolution analysis
PnP-MUNIT relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (i.e. the spatial resolution of the content relative to the images) to explain this effect. We simulated 4 datasets, denoted as RefRes-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution. For each dataset, we trained 3 content/style models with content capacity denoted as ContentRes-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of downsampling blocks in the networks. E.g. ContentRes-2 model produced content maps half the spatial resolution of the ContentRes-1 model.
Figure 5: The effect of misaligned reference on the reconstruction and the robustness of the CR module to it.
Figure 6: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the $\gamma$ parameter. $\gamma$ controls the consistency of the content with the measured k-space.
Figure 7: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices.
Figure 8: Two examples from the NYU benchmark. (a) Comparison of all the methods at the highest acceleration of $R$=5. (b) Comparison between PnP-MUNIT and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
Table IV compares PnP-MUNIT reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-2 and RefRes-4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, Refsize-2 and RefRes-4 datasets had similar levels of shared structure despite the lower reference resolution in the latter, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-8 case, PnP-MUNIT dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-MUNIT to be effective, specifically 1/42=1/16 times the local information of the full-resolution case.
V-A3 Convergence
We explored the convergence of PnP-MUNIT by comparing it with two variants. The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT. The second used reference content $\hat{c}$ with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the ContentRes-1 model additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 3 shows the convergence curves and intermediate reconstructions. The upper-bound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
V-A4 Effectiveness in resolving contrast-specific structure
@@
V-A5 Sensitivity to misalignment of the reference
While PnP-MUNIT expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2 rotation in the reference image and performed reconstruction with and without the CR module, comparing also with PnP-MUNIT that used an aligned reference image. As shown in Fig. 5, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.

### v3

@@ 4.1.2 Benchmark on the NYU DICOM dataset
Table 2: LUMC data split. The splits were made at the subject-level, i.e. each subject belonged to exactly one subset. Splits containing only image-domain data are marked with I. The recon-val and recon-test splits additionally include raw T2W multi-coil k-space and coil sensitivities.
@@
4.1.3 Benchmark and ablation on the LUMC multi-coil dataset
@@
In our benchmark on clinical multi-coil data, the goal was to test PnP-CoSMo on a constrained real-world problem where only the image-domain data was available for training. This is often the case, as in clinical practice raw data is discarded after acquisition and only the final reconstructions are retained. Moreover, the T1W/T2W images were not fully subject-wise paired, representing a realistic case of data imbalance. Our in-house data consisted of brain scans of patients from LUMC, the use of which was approved for research purpose by the institutional review board. A total of 1669 brain scans were obtained from 817 clinical MR examinations of 630 patients acquired on 3T Philips Ingenia scanners. We focused on accelerating two T2W sequences – (a) 2D T2W TSE transversal and (b) 2D T2W TSE coronal. For guidance, two corresponding T1W sequences were used – (a) 3D T1W TFE transversal and (b) 3D T1W TSE coronal. The transversal and coronal protocols were considered as two separate datasets, namely LUMC-TRA and LUMC-COR. Table 1 shows an overview of the four sequences. Note that the in-plane resolution of the T1W scans was 1.3-2.5 times as low as that of the T2W scans, making these datasets more challenging for T1W-guided T2W reconstruction.
@@
As with the NYU dataset, the LUMC datasets were split into model-train, model-val, recon-val, and recon-test. The former two splits contained only image-domain data, which was used to train and validate the content/style models. The models were pre-trained on the full model-train splits, ignoring any pairing between T1W and T2W scans. Additionally, 20 subjects from the model-train split were designated for PFT where the pairing information was used in training and the reference scans were aligned via registration. The recon-val and recon-test splits additionally included multi-coil T2W raw data, which comprised 6-channel (LUMC-TRA) and 13-channel (LUMC-COR) k-space and coil sensitivity maps. This k-space was already undersampled (1D Cartesian random) at acquisition with clinical acceleration of $R$=1.8-2. We further undersampled it retrospectively to higher accelerations $R\in\{4,6,8,10\}$ by dropping subsets of the acquired lines. An overview of the data split is shown in Table 2. Spatially aligned images required by all but the unpaired training set were obtained via rigid registration and the reference T1W images were resampled to the T2W resolution. During unpaired training, all images were resampled to the median T2W resolution.
@@
Given the of absence of k-space training data, end-to-end reconstruction methods were not feasible as baselines. Hence, we compared PnP-CoSMo with only the feasible types of baselines, i.e. classical, semi-classical, and plug-and-play reconstruction and image-to-image translation. Among classical methods, we used the unguided L1-wavelet CS (CS-WT) and the guided STV-based CS (CS-STV) (Ehrhardt and Betcke, 2016). As an unguided plug-and-play baseline, we used PnP-CNN (Ahmad et al., 2020), and as a representative image translation baseline, we compared against MUNIT itself. Using MUNIT, deterministic image translation was approximated by combining the reference content with 200 randomly sampled T2W style codes and taking a pixel-wise mean of the resulting synthetic images to obtain a single synthetic image. As semi-classical guided baseline, we used (PROSIT) (Mattern et al., 2020), which combines deterministic image translation with L2-regularized least-squares reconstruction. Additionally, as an ablation study for PnP-CoSMo, we ablated PFT and CR to assess their contribution, and finally, as an upperbound for PnP-CoSMo representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c^{*}$. In both NYU and LUMC benchmarks, we used three perceptual metrics for evaluation – SSIM, HaarPSI, and DISTS. While SSIM is used commonly, HaarPSI and DISTS are known to correlate better with visual judgment of image quality (Kastryulin et al., 2023). All three metrics are bounded in $[0,1]$ where $1$ represents perfect image quality. We conducted paired Wilcoxon signed-rank tests to measure statistical significance when comparing pairs of algorithms.
4.1.4 Radiological evaluation
@@
Finally, as an extension to the LUMC benchmark, we conducted a radiological evaluation assessing the visual and diagnostic quality of PnP-CoSMo reconstructions. The study was conducted on a small sample of the LUMC-TRA recon-test set. With the help of a junior radiologist, we selected 3 cases with brain metastases. The evaluation was performed by a senior neuroradiologist and comprised two sets of criteria, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. A diagnostic-quality reconstruction was defined as one that scored at least ”fairly diagnostic” on the overall diagnostic quality criterion. We evaluated PnP-CoSMo and PnP-CNN at 4 clinically realistic accelerations of $R\in\{3,4,5,6\}$. In total, 9 images per patient, including the (4$\times$2=8) reconstructions and the clinical ground truth (acquired at $R$=2 and reconstructed by the vendor software), were presented to the senior neuroradiologist who scored each image individually, blinded to the reconstruction method and $R$.
4.2 Implementation details
@@
We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of specifies number of channels. We used 2 or 4 channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal. To stabilize GAN training, we used spectral normalization (Miyato et al., 2018). For hyperparameters, we found that pre-training loss weight values $\alpha_{1}=\alpha_{2}=\alpha_{3}=1$ and paired fine-tuning values $\beta_{1}=\beta_{2}=\beta_{3}=1$ worked generally well for the in vivo datasets. In the benchmarks, the step size $\gamma$ of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
@@
In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1. In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations. All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009). All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
5 Results
5.1 Experiments on simulated MR datasets
5.1.1 Perturbation of disentanglement strength
PnP-CoSMo assumes that the content and style aspects of the images are sufficiently disentangled. Here, we evaluated the effect of perturbing the disentanglement loss weights $\alpha_{2}$ and $\alpha_{3}$ (Eq. (6)) of MUNIT on the PnP-CoSMo reconstruction quality. Table 3 shows the reconstruction PSNR for different weight values. We observed that optimal disentanglement, and thus the best reconstruction, was achieved at $\alpha_{2}=\alpha_{3}=0.1$. At lower values, poor disentanglement of content and style in the model resulted in poor PnP-CoSMo reconstruction quality. This was expected since PnP-CoSMo relies on the assumption that the content representation is (sufficiently) contrast-independent, which was less enforced at these levels. At higher values, the training over-emphasized disentanglement while under-emphasizing the GAN and image recovery loss terms, thereby leading to worse preservation of image information in the latent space and thus, to worse reconstructions.
@@ 5.1.1 Perturbation of disentanglement strength
Table 3: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices). PSNR worse than the conventional L1-wavelet CS reconstruction is marked as .
@@
Table 4: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal content capacity is shown in bold. PSNR worse than the conventional L1-wavelet CS reconstruction is marked as .
@@
RefRes-1 RefRes-1/2 RefRes-1/4 RefRes-1/8
$J_{M}$=1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
$J_{M}$=(1/2)2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
$J_{M}$=(1/4)2 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
5.1.2 Content capacity analysis
PnP-CoSMo relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (Eq. (14)) to explain this effect. We simulated 4 datasets, denoted as RefRes-$1/n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, the T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components, while maintaining the same spatial resolution. For each dataset, we trained 3 content/style models with content capacity $J_{M}$=$(1/m)^{2}$, where $m\in\{1,2,4\}$ is the content downsampling factor, which depends on the number of up/downsampling blocks in the networks. E.g. the $J_{M}$=(1/2)2 model produced content maps half the spatial resolution of the $J_{M}$=1 model.
Figure 4: (a) Convergence of PnP-CoSMo and its two variants at $R$=2. (b) Evolution of the reconstruction shown with the corresponding error map.
Table 4 compares PnP-CoSMo reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below the conventional L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-1/2 and RefRes-1/4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, RefRes-1/2 and RefRes-1/4 datasets had similar levels of shared structure (i.e. $J_{M}$=1/22=1/4 times the full-resolition case) despite the lower reference resolution in RefRes-1/4, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-1/8 case, PnP-CoSMo dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-CoSMo to be effective, specifically $J_{M}$=1/42=1/16 times the local information of the full-resolution case.
5.1.3 Convergence
We explored the convergence of PnP-CoSMo by comparing it with two variants. The first used true content $c^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upperbound of PnP-CoSMo. The second used reference content $\hat{c}$ with CR module disabled, representing a lowerbound of PnP-CoSMo where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the $J_{M}$=1 model which was additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 4 shows the convergence curves and intermediate reconstructions. The upperbound version, given $c^{*}$, converged in a single iteration, while with $\hat{c}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR module closed the gap with the upperbound, although at the cost of slower convergence. It is also observed that the content discrepancy reduced when using CR and that the style estimate always converged within the first few iterations, as we hypothesized in Section 3.3.
5.1.4 Effectiveness in resolving contrast-specific structure
@@
5.1.5 Sensitivity to misalignment of the reference
Figure 6: A misaligned reference image severely affects the reconstruction when CR is disabled. Enabling CR makes the reconstruction process robust to small misalignments.
While PnP-CoSMo expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2 rotation in the reference image and performed reconstruction at $R$=4 with and without the CR module, comparing also with PnP-CoSMo that used an aligned reference image. As shown in Fig. 6, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
5.1.6 Influence of the CR step size
Figure 7: Reconstruction PSNR as a function of acceleration $R$ and k-space noise level $\sigma$ for different values of the hyperparameter $\gamma$. $\gamma$ is the CR step size controlling the consistency of the content with the measured k-space.
The hyperparameter $\gamma$ in Algorithm 1 controls the CR step size. Fig. 7 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreased with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
5.2 Benchmark on the NYU DICOM data
In the NYU benchmarking setup considered here, let the T1W-guided T2W reconstruction task be denoted as Task-1 and the T2W-guided T1W reconstruction task as Task-2. Fig. 8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two tasks. We observe three main trends. First, the end-to-end methods generally dropped in performance in the out-of-distribution cases ($p$$<$0.05 throughout, with the exception of SSIM at $R$=5 and HaarPsi at $R\in${2,4,5} for MoDL on Task-2). Second, PnP-CoSMo outperformed out-of-distribution models ($p$$<$0.05, except DISTS of MoDL and MCVarNet at $R$=2 on Task-2). Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind. Third, considering the in-distribution model performance, PnP-CoSMo largely outperformed MoDL ($p$$<$0.05 throughout, except SSIM at $R$=2 and DISTS at $R\in${2,3,4} on Task-1 and DISTS at $R$=2 on Task-2) and MTrans ($p$$<$0.05 throughout, except SSIM at $R\in${3,4} and DISTS at $R\in${2,3}), and was roughly comparable to MC-VarNet, despite having lenient training data requirements. As shown in Fig. 9a, PnP-CoSMo could resolve certain fine details even at the high acceleration of $R$=5, which the baselines failed at. In the pathological case shown in Fig. 9b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite lower metrics, when many features of edema were absent in the reference image.
Figure 8: Evaluation plots over recon-test subset of NYU DICOM dataset consisting of 550 slices for the two reconstruction tasks. ID and OOD refer to in-distribution and out-of-distribution models, respectively. Note that MoDL is a single-contrast baseline and hence did not use any reference images. Unlike other methods, PnP-CoSMo, by design, required only a single content/style model for both tasks.
Figure 9: Two examples from the NYU benchmark for Task-1 (i.e. T1W-guided T2W reconstruction). (a) Comparison of PnP-CoSMo with the in-distribution end-to-end models at the highest acceleration of $R$=5. (b) Comparison between PnP-CoSMo and MC-VarNet at $R$=2 and $R$=4 on a pathological case where many features of the edema are T2W contrast-specific.
5.3 Benchmark and ablation on the LUMC multi-coil data